OpenStudy (anonymous):

2n-7 = 15 ---- 3 @sammixboo

sammixboo (sammixboo):

Does the problem look like this? 2n - \(\rm \large \frac{7}{3}\) = 15?

sammixboo (sammixboo):

Or \(\rm \large \frac{2n~-~7}{3}\) = 15

sammixboo (sammixboo):

I will chose my second choice =) Either way they end up being the same answer.

sammixboo (sammixboo):

So first you must multiply the 3 on both sides of the equation \(\rm \large \frac{2n~-~7}{3} * 3\) = 15 * 3 2n - 7 = 15 * 3 So what is 15*3?

OpenStudy (shadowlegendx):

Inverse Operations c;

OpenStudy (shadowlegendx):

@sammixboo I know this one c;

sammixboo (sammixboo):

;) Puhlease I got this

sammixboo (sammixboo):

Go away, Batman

sammixboo (sammixboo):

=)

sammixboo (sammixboo):

I miss your old profile picture :(

sammixboo (sammixboo):

I didn't know O:)

sammixboo (sammixboo):

Ppor guy.. We are spamming his post

OpenStudy (anonymous):

BATMAN!!

sammixboo (sammixboo):

OKAAAY well since he is not responding 15 * 3 = 45, so now we plug is 45 were 15 * 3 was. 2n - 7 = 45 Next we have to add 7 on both sides, since we are subtracting 7 in the original problem 2n - 7 + 7 = 45 + 7 2n = 45 + 7

sammixboo (sammixboo):

So what is 45 + 7

sammixboo (sammixboo):

Why thank you, Batman

OpenStudy (anonymous):

52

sammixboo (sammixboo):

Yes~ Now we plug in 52 where 45 + 7 was 2n = 52

sammixboo (sammixboo):

Do you know what we do nex?

OpenStudy (anonymous):

no what do we do :D

sammixboo (sammixboo):

We multiply 2 on both sides So what is 52 x 2?

OpenStudy (anonymous):

104

sammixboo (sammixboo):

Yes so n = 104

OpenStudy (anonymous):

yay